OpenStudy (muhammadturawa):

Rationalise the Numerator: (√4+h - 2)/h

OpenStudy (anonymous):

\[\frac{\sqrt{4+h}-2}{h}\times \frac{\sqrt{4+h}+2}{\sqrt{4+h}+2}\] is a start

OpenStudy (anonymous):

leave the denominator in factored form the numerator uses \((a-b)(a+b)=a^2-b^2\) to get \[4+h-4=h\]